OpenStudy (anonymous):

What is the slope of the line that passes through the points (-4,-3) and (10, 4)?

OpenStudy (jaredstone4):

m (slope) = (y2 - y1)/(x2-x1) m = (4-(-3))/(10-(-4)) m = 7/14 m = 1/2

OpenStudy (jdoe0001):

http://mathb.in/12793

OpenStudy (amorfide):

To get the y value, would be to get the length of the line going vertical. To do that you must do the highest y value, subtract the lower y value, to get the length. to get the x value, would be to get the length of the line going horizontal. Do the same, highest x subtract the lowest x value. the slope is equal to y/x